Fifteen pounds of trail mix are made by mixing 3 pounds of raisins worth $3 per pound with peanuts worth $4 per pound and M&Ms worth $5 per pound. How many pounds of peanuts must be used if the mixture is to be worth $4.20 per pound?
Answer by richwmiller(17219)

3*3+4x+5y=15*4.20
x+y=12
x=6 pounds of peanuts and y=6 pounds of M&M's
check
3*3+4*6+5*6=15*4.20
ok